OpenStudy (zubhanwc3):

find the particular solution of the differential equation that satisfies the initial condition

OpenStudy (zubhanwc3):

\[\sqrt{x}+\sqrt{y}y'=0\]\[y(1) = 4\]

OpenStudy (zubhanwc3):

@zepdrix

zepdrix (zepdrix):

So this is separable, yes?

zepdrix (zepdrix):

Subtract sqrt x from each side, "multiply" the dx over,\[\Large\rm \sqrt y ~dy=-\sqrt x~ dx\]

OpenStudy (zubhanwc3):

these were 2 equations that were in the same line for the problem

zepdrix (zepdrix):

Understand how to integrate each side?

OpenStudy (zubhanwc3):

i have a general idea, but could you go over it, im in online school, and my calc foundation is very weak :/

zepdrix (zepdrix):

\[\Large\rm \int\limits\sqrt y ~dy=\int\limits -\sqrt x~ dx\]So you want to `rationalize` the terms before you integrate.\[\Large\rm \int\limits y^{1/2} ~dy=\int\limits -x^{1/2}~ dx\]Remember that from your rules of exponents I hope? :O Then we just apply the Power Rule for Integration to each term.\[\Large\rm \color{royalblue}{\int\limits x^{n}~dx=\frac{1}{n+1}x^{n+1}}\]

OpenStudy (zubhanwc3):

thank you, i know what to do from here.

zepdrix (zepdrix):

cool c: